OpenStudy (alex6799):

Find the exact value. If the expression is undefined, write undefined. csc 135 A. B. 0 C. undefined D.

OpenStudy (alex6799):

\[A. \sqrt{2}\] \[D. \frac{ 1 }{ 2 }\]

OpenStudy (aihberkhan):

What is your expression?

OpenStudy (aihberkhan):

Oh its csc 135?

OpenStudy (alex6799):

yea

OpenStudy (aihberkhan):

Okay. So if you were to look at a reference angle. \(135º\) is in the 4th quadrant with a of \(45º\). Now... \(sin 45º\) is \(√2 /2\). This means that \(csc 45º = √2 \) Since, this is positive and it is in the 5th quadrant... the answer would be A! \(√2 \)

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @alex6799

OpenStudy (alex6799):

YAY I GOT THIS ONE RIGHT!!!!!!